# AffordDexGrasp: Open-set Language-guided Dexterous Grasp with Generalizable-Instructive Affordance

### v1

@@ Abstract
Language-guided robot dexterous generation enables robots to grasp and manipulate objects based on human commands. However, previous data-driven methods are hard to understand intention and execute grasping with unseen categories in the open set. In this work, we explore a new task, Open-set Language-guided Dexterous Grasp, and find that the main challenge is the huge gap between high-level human language semantics and low-level robot actions. To solve this problem, we propose an Affordance Dexterous Grasp (AffordDexGrasp) framework, with the insight of bridging the gap with a new generalizable-instructive affordance representation. This affordance can generalize to unseen categories by leveraging the object’s local structure and category-agnostic semantic attributes, thereby effectively guiding dexterous grasp generation. Built upon the affordance, our framework introduces Affordacne Flow Matching (AFM) for affordance generation with language as input, and Grasp Flow Matching (GFM) for generating dexterous grasp with affordance as input. To evaluate our framework, we build an open-set table-top language-guided dexterous grasp dataset. Extensive experiments in the simulation and real worlds show that our framework surpasses all previous methods in open-set generalization.
@@ 1 Introduction
Recent works explore the task of language-guided dexterous grasp generation[51, 20, 21, 59], aiming to enable dexterous hands to perform actions based on language instructions, going beyond previous works that focus on stable grasping [46, 54, 26, 38]. The typical language-guided approaches employ language as the condition of generative models to predict hand parameters, achieving impressive performance on objects within known categories [51, 4]. However, in the open real world, there are many categories that may not appear during training, and the cost of collecting data for dexterous hands is quite expensive. Therefore, the open-set generalization on unseen category samples is crucial for robot grasp. While previous works on parallel grippers explore open-set tasks [36, 23], there is limited research on open-set generalization for dexterous hands, as their significantly higher degrees of freedom present complex challenges.
@@
To solve the above challenge, we propose the Affordance Dexterous Grasp (AffordDexGrasp) framework, with the key insight of employing a new affordance as an intermediate representation to bridge the gap between high-level language and low-level grasp actions. Our affordance representation combines two characteristics: generalizable (generalizing to unseen categories based on language) and instructive (guiding dexterous grasp generation effectively). However, achieving these two characteristics is not trivial. For example, as shown in Figure 2, fine-grained contact information can effectively guide grasp generation or optimization [2, 9, 17, 55], but it is difficult to generalize to unseen categories. In contrast, coarse information, such as object parts, can be obtained from a pre-trained computer vision model [27, 43, 41], but it is too coarse to guide dexterous hand actions with higher degrees of freedom.
To achieve these characteristics, we propose the Generalizable-Instructive Affordance by defining a general graspable affordance that aligns with category-independent information, such as intention, object parts, and direction. As shown in Figure 2, the affordance represents the potential graspable regions of all grasps with the same semantics. In this way, the models do not need to learn complex dexterous contact patterns but instead focus on a general graspable area, which can be well aligned with category-agnostic semantic attributes and guide grasp generation effectively.
@@ 2.1 Dexterous Grasp
Dexterous grasping is critical in robotics, which equips robots with human-like grasping capabilities. Some methods [29, 58, 44, 48] focus on grasp stability and quality, while recent studies explore task-oriented [50, 61] or language-guided grasping [21, 51] with specific semantic intention. For language-guided task, the data-driven methods achieve impressive performance in unseen object within seen categories by leveraging language conditioned generative model [4, 51]. However, we find that these models are hard to generalize for samples from unseen categories. And we find that the main challenge is the huge gap between language and grasp action. To solve this problems, we propose AffordDexGrasp framework with the insight that bridge the gap by generalizable-instructive affordance.
Figure 2: Different affordance representations. (a) While conatct map are too elaborate to generalize and object part are too coarse to guide grasping, our affrodacne achieve a balance. (b) While object part has a lower upper bound and contact shows significant degradation in generalization, only our affordance effectively achieve the balance (Top-1 indicates grasp intention consistency).
@@ 2.2 Open-set Robot Grasp
Exploring the performance of robotic models in open-set scenarios is crucial due to the diverse object categories in the real world and the high cost of data collection. For tasks that only focus on stable grasping, both parallel grippers [6, 52] and dexterous hands [47, 58] achieve impressive performance on open set. However, when it comes to considering task-oriented [45] or language-guided grasping [33], it becomes much more difficult. Some works in parallel grippers achieve this by using pre-trained visual grounding models [22, 41, 27], implicit feature semantic fields [18, 36, 32], or multimodal large models [13, 14]. Compared to parallel grippers, dexterous hands have a significantly higher number of degrees of freedom, making the methods of parallel grippers difficult to transfer to dexterous hand [51]. In this paper, we explore the open-set language-guided dexterous grasping task and propose a novel framework to address it.
@@ 2.3 Grasp Affordance
Affordance is first introduced in [8], referring to environmental action possibilities. There are currently two types of grasp affordance: fine-grained contact [42, 3] and coarse-grained object segmentation [27]. The fine-grained information, such as contact areas, normals, or key points, is commonly used in hand-object interaction[59, 62, 9, 2]. However, this fine-grained information is difficult to generalize well to unseen categories. As a result, inaccurate contact information would lead to unreasonable grasping. On the other hand, We find through experiments that the coarse object segmentation [27] is too coarse to guide dexterous grasp generation with relatively high degrees of freedom. To solve this problem, we propose Generalizable-Instructive Affordance, which can generalize through the local structure of objects using category-independent clues and effectively guide grasp generation.
@@ 3 Affordance Dexterous Grasp Framework
Given the scene point clouds $\mathcal{O}$, RGB images $\mathcal{I}$ and language guidance $\mathcal{L}$ as input, our goal is to generate intention aligned and high-quality dexterous grasps poses $\mathcal{G}^{dex}$ = $(r,t,q)$, where $r$ denotes the rotation, $t$ denotes the translation, and $q$ represents the joint angles of the dexterous hand.
@@ 3.2.2 Affordance Flow Matching
Specifically, the input scene object point cloud $\mathcal{O}$ is encoded by Pointnet++ [34], language is encoded by a pre-trained CLIP model [35], and the direction vector and time embedding are encoded by the Multilayer Perceptrons (MLP). Due to the permutation invariance of point level affordance, the noisy affordance are concatenated with object features to aware the position and sematic information. Then all features are fed into a Perceiver IO Transformer, following [16, 49]. The noisy affordance features are served as input and output array, and the features of language and direction are served as latent array. The outputs of the decoder go through an MLP to obtain the prediction of the flow $v_{\eta}(\Omega_{t},t|c)\in R^{M}$. Finally, in the inference time, we adopt mulitple steps $1/\Delta t$ following the equation:
@@ 3.2.3 Grasp Flow Matching
We introduce Grasp Flow Matching (GFM) to learn the grasp distribution, conditioned on the output of MLLM and AFM. GFM is also built upon the flow matching model, and the objective is to generate a parameterized dexterous hand pose $\mathcal{G}^{dex}$ = $(r,t,q)$, which represents the global rotation, translation, and joint angles. In the training time, grasp generation requires calculating explicit losses in 3D space, which involves obtaining the target dexterous poses, which go through forward kinematics to obtain hand configuration [53]. To obtain the target pose, we estimate it in one step during the training stage:
@@ 4 Open-set Dexterous Grasp Dataset
- Open Set Data Split. We divide the categories in our dataset into two groups, A and B, excluding one from the training set each time to form Open Set A and B. Open Set A consists of 9,688 samples of 808 objects from 10 unseen categories, and 159 unseen objects from seen categories. Open Set B contains 29,744 samples of 568 unseen objects from 10 unseen categories, and 202 unseen objects from seen categories.
@@ 5.2 Implementation Details
For training our framework, the number of epochs is set to 50 for AFM and 200 for GFM. For GFM, the loss weight of L2 loss is set to 1. For GFM, we set $\lambda_{\text{pose}}=10$, $\lambda_{\text{chamfer}}=1$, and $\lambda_{\text{tip}}=2$. The batch size is set to 16 for AFM and 64 for GFM. The initial learning rate is $2.0\times 10^{-4}$, decaying to $2.0\times 10^{-5}$ using a cosine learning rate scheduler [31]. The Adam optimizer is used with a weight decay rate of $5.0\times 10^{-6}$. In the inference, the time step is set to 10 for AFM and 20 for GFM. For the optimization, $\lambda_{1}=100$, $\lambda_{2}=10$, $\lambda_{3}=100$, $\lambda_{4}=10$, $\lambda_{5}=100$. The number of optimization iterations is set to 200. All experiments are implemented using PyTorch on a single RTX 4090 GPU.
@@ 5.3 Comparison with SOTA Methods
We compare our methods with the SOTA methods, as shown in Table 1. The generic grasp methods are reproduced by concatenating the point cloud features and features of same language guidance with ours. The same encoders are employed for fair comparison, and the penetration loss is excluded to avoid learning difficulties according to [51]. The results show that our framework significantly outperforms all previous methods in terms of intention consistency and grasp quality. Our framework also achieves a reasonably high level of diversity, as excessive diversity may lead to unnatural postures. Previous language-guided methods fail to generalize well to unseen categories due to the huge gap between language and grasping actions. Similarly, contact-based methods don’t perform well, as the contact maps are difficult to generalize. Additionally, our method outperforms other methods in the close set, as shown in Table 3. Overall, the results indicate that our framework achieves strong performance in intention-aligned, high-quality, and diverse grasps.
@@ 5.6 Real World Experiments
The real-word experiments are conducted to verify the simulation-to-reality ability of our framework. We employ a Leap Hand, a Kinova Gen3 6DOF arms and an original wrist RGB-D camera of Kinova arm. And we collect several common objects in daily life, including unseen objects and unseen categories. In experiment, we synthesize the scene point cloud by taking several partial depth maps around the object. Then the scene point cloud, a RGB image and the user language command are fed into our framework to obtain the dexterous grasp pose. During execution, we first move the the arm to a pre-grasp position, then synchronously move the joints of the robotic arm and the dexterous hand to reach the target pose. In evaluation, a grasp is considered as success if the hand can lift it up and the action is consisted with the given command. The results shown in Table 5 shows that our framework achieve higher success rate than previous method.

### v2

@@ Abstract
Language-guided robot dexterous generation enables robots to grasp and manipulate objects based on human commands. However, previous data-driven methods are hard to understand intention and execute grasping with unseen categories in the open set. In this work, we explore a new task, Open-set Language-guided Dexterous Grasp, and find that the main challenge is the huge gap between high-level human language semantics and low-level robot actions. To solve this problem, we propose an Affordance Dexterous Grasp (AffordDexGrasp) framework, with the insight of bridging the gap with a new generalizable-instructive affordance representation. This affordance can generalize to unseen categories by leveraging the object’s local structure and category-agnostic semantic attributes, thereby effectively guiding dexterous grasp generation. Built upon the affordance, our framework introduces Affordance Flow Matching (AFM) for affordance generation with language as input, and Grasp Flow Matching (GFM) for generating dexterous grasp with affordance as input. To evaluate our framework, we build an open-set table-top language-guided dexterous grasp dataset. Extensive experiments in the simulation and real worlds show that our framework surpasses all previous methods in open-set generalization.
@@ 1 Introduction
Recent works explore the task of language-guided dexterous grasp generation[24, 25, 69, 20], aiming to enable dexterous hands to perform actions based on language instructions, going beyond previous works that focus on stable grasping [54, 31, 44]. The typical language-guided approaches employ language as the condition of generative models to predict hand parameters, achieving impressive performance on objects within known categories [59, 4]. However, in the open real world, there are many categories that may not appear during training, and the cost of collecting data for dexterous hands is quite expensive. Therefore, the open-set generalization on unseen category samples is crucial for robot grasp. While previous works on parallel grippers explore open-set tasks [42, 27, 70], there is limited research on open-set generalization for dexterous hands, as their significantly higher degrees of freedom present complex challenges.
@@
To solve the above challenges, we propose the Affordance Dexterous Grasp (AffordDexGrasp) framework, with the insight of employing a new affordance as an intermediate representation to bridge the gap between high-level language and low-level grasp actions. Our affordance representation combines two characteristics: generalizable (generalizing to unseen categories based on language) and instructive (guiding dexterous grasp generation effectively). However, achieving these two characteristics is not trivial. For example, as shown in Figure 2, fine-grained contact information can effectively guide grasp generation or optimization [2, 9, 21, 64, 19], but it is difficult to generalize to unseen categories. In contrast, coarse information, such as object parts, can be obtained from a off-the-shelf model [32, 50, 48], but it is too coarse to guide dexterous hand actions with higher degrees of freedom.
To achieve these characteristics, we propose the Generalizable-Instructive Affordance by defining a general dexterous affordance that aligns with category-independent information, such as intention, object parts, and direction. As shown in Figure 2, the affordance represents the potential graspable regions of all grasps with the same semantics. In this way, the models do not need to learn complex dexterous contact patterns but instead focus on a general graspable area, which can be well aligned with category-agnostic semantic attributes and guide grasp generation effectively.
@@ 2.1 Dexterous Grasp
Dexterous grasping is critical in robotics, which equips robots with human-like grasping capabilities [61, 29]. Some methods [34, 68, 51, 56] focus on grasp stability and quality, while recent studies explore task-oriented [58, 72] or language-guided grasping [25, 59] with specific semantic intention. For language-guided task, the data-driven methods achieve impressive performance in unseen object within seen categories by leveraging language conditioned generative model [4, 59]. However, we find that these models are hard to generalize for samples from unseen categories. And we find that the main challenge is the huge gap between language and grasp action. To solve this problems, we propose AffordDexGrasp framework with the insight that bridge the gap by generalizable-instructive affordance.
Figure 2: Different affordance representations. (a) While contact map are too elaborate to generalize and object part are too coarse to guide grasping, our affrodacne achieve a balance. (b) While object part has a lower upper bound and contact shows significant degradation in generalization, only our affordance effectively achieve the balance (Top-1 indicates grasp intention consistency).
@@ 2.2 Open-set Robot Grasp
Exploring the performance of robotic models in open-set scenarios is crucial due to the diverse object categories in the real world and the high cost of data collection. For tasks that only focus on stable grasping, both parallel grippers [6, 60, 52] and dexterous hands [55, 68] achieve impressive performance on open set. However, when it comes to considering task-oriented [53] or language-guided grasping [38], it becomes much more difficult. Some works in parallel grippers achieve this by using pre-trained visual grounding models [26, 48, 32], implicit feature semantic fields [22, 42, 37], or multimodal large models [14, 15]. Compared to parallel grippers, dexterous hands have a significantly higher number of degrees of freedom, making the methods of parallel grippers difficult to transfer to dexterous hand [59]. In this paper, we explore the open-set language-guided dexterous grasping task and propose a novel framework to address it.
@@ 2.3 Grasp Affordance
Affordance is first introduced in [8], referring to environmental action possibilities. There are currently two types of grasp affordance: fine-grained contact [49, 3] and coarse-grained object segmentation [32]. The fine-grained information, such as contact areas, normals, or key points, is commonly used in hand-object interaction[69, 73, 9, 2, 66]. However, this fine-grained information is difficult to generalize well to unseen categories. As a result, inaccurate contact information would lead to unreasonable grasping. On the other hand, We find through experiments that the coarse object segmentation [32] is too coarse to guide dexterous grasp generation with relatively high degrees of freedom. To solve this problem, we propose Generalizable-Instructive Affordance, which can generalize through the local structure of objects using category-independent clues and effectively guide grasp generation.
@@ 3 Affordance Dexterous Grasp Framework
[…]
@@ 3.2.2 Affordance Flow Matching
Specifically, the input scene object point cloud $\mathcal{O}$ is encoded by Pointnet++ [40], language is encoded by a pre-trained CLIP model [41], and the direction vector and time embedding are encoded by the Multilayer Perceptrons (MLP). Due to the permutation invariance of point level affordance, the noisy affordance are concatenated with object features to aware the position and semantic information. Then all features are fed into a Perceiver IO Transformer, following [17, 57]. The noisy affordance features are served as input and output array, and the features of language and direction are served as latent array. The outputs of the decoder go through an MLP to obtain the prediction of the flow $v_{\eta}(\Omega_{t},t|c)\in R^{M}$. Finally, in the inference time, we adopt mulitple steps $1/\Delta t$ following the equation:
@@ 3.2.3 Grasp Flow Matching
[…]
@@ 4 Open-set Dexterous Grasp Dataset
- Open Set Data Split. To enable comprehensive evaluation, we construct Open sets A and B using two independent dataset splits, each with its own set of unseen categories. For each split, all categories are first labeled as either seen or unseen. Then, 80% of the objects from the seen categories are used for training, while the remaining 20% and all objects from the unseen categories are used for testing.
@@ 5.2 Implementation Details
For training our framework, the number of epochs is set to 50 for AFM and 200 for GFM. For AFM, the loss weight of L2 loss is set to 1. For GFM, we set $\lambda_{\text{pose}}=10$, $\lambda_{\text{chamfer}}=1$, and $\lambda_{\text{tip}}=2$. The batch size is set to 16 for AFM and 64 for GFM. The initial learning rate is $2.0\times 10^{-4}$, decaying to $2.0\times 10^{-5}$ using a cosine learning rate scheduler [36]. The Adam optimizer is used with a weight decay rate of $5.0\times 10^{-6}$. In the inference, the time step is set to 10 for AFM and 20 for GFM. For the optimization, $\lambda_{1}=100$, $\lambda_{2}=10$, $\lambda_{3}=100$, $\lambda_{4}=10$, $\lambda_{5}=100$. The number of optimization iterations is set to 200. All experiments are implemented using PyTorch on a single RTX 4090 GPU.
@@ 5.3 Comparison with SOTA Methods
We compare our methods with the SOTA methods, as shown in Table 1. The generic grasp methods are reproduced by concatenating the point cloud features and features of same language guidance with ours. The same encoders are employed for fair comparison, and the penetration loss is excluded to avoid learning difficulties according to [59]. The results show that our framework significantly outperforms all previous methods in terms of intention consistency and grasp quality. Our framework also achieves a reasonably high level of diversity, as excessive diversity may lead to unnatural postures. Previous language-guided methods fail to generalize well to unseen categories due to the huge gap between language and grasping actions. Similarly, contact-based methods don’t perform well, as the contact maps are difficult to generalize. Additionally, our method outperforms other methods in the close set, as shown in Table 3. Overall, the results indicate that our framework achieves strong performance in generating intention-aligned, high-quality, and diverse grasps.
@@ 5.6 Real World Experiments
The real-word experiments are conducted to verify the simulation-to-reality ability of our framework. We employ a Leap Hand, a Kinova Gen3 6DOF arms and an original wrist RGB-D camera of Kinova arm. And we collect several common objects in daily life, including unseen objects and unseen categories. In experiment, we synthesize the scene point cloud by taking several partial depth maps around the object. Then the scene point cloud, a RGB image and the user language command are fed into our framework to obtain the dexterous grasp pose. During execution, we first move the the arm to a pre-grasp position, then synchronously move the joints of the robotic arm and the dexterous hand to reach the target pose. In evaluation, a grasp is considered as success if the hand can lift it up and the action is consistent with the given command. The results shown in Table 5 shows that our framework achieve higher success rate than previous method.